OpenStudy (loser66):

Converge or diverge? \[\sum_{n=0}^\infty \dfrac{\sqrt{n+1}}{n^2-3n+1}\] Please, help

OpenStudy (loser66):

@rational

OpenStudy (rational):

are you allowed to use comparison test ?

OpenStudy (loser66):

Yes

OpenStudy (irishboy123):

.

OpenStudy (rational):

For \(n\gt 4\) we have \[\dfrac{\sqrt{n+1}}{n^2-3n+1}\lt \dfrac{\sqrt{n+1}}{n^2-3n-4}=\dfrac{\sqrt{n+1}}{(n+1)(n-4)}\\=\dfrac{1}{\sqrt{n+1}(n-4)}\lt \dfrac{1}{\sqrt{n-4}(n-4)}=\dfrac{1}{(n-4)^{3/2}}\] Since \(\frac{3}{2}\gt 1\), the series converges by p-series test.

OpenStudy (loser66):

If it is so, we depend on n is not 4, right? while the original one is not. How to argue? because original one gave us the rank from 0 to infinitive

OpenStudy (rational):

Easy, split the series into two : \[\sum_{n=0}^\infty \dfrac{\sqrt{n+1}}{n^2-3n+1}=\sum_{n=0}^4 \dfrac{\sqrt{n+1}}{n^2-3n+1}+\sum_{n=5}^\infty \dfrac{\sqrt{n+1}}{n^2-3n+1}\] the first series on right hand side evaluates to some number, and the second series converges by comparison with p-series. so the overall series converges.

OpenStudy (loser66):

I got it. Thank you so much. Is there any way for my second one?

OpenStudy (rational):

is there a typo in second series the index variable "n" doesn't agree with the varibale in expression "k"

OpenStudy (loser66):

I am sorry, it is k =1 to infinitive

OpenStudy (loser66):

\[\sum_{k=1}^\infty \dfrac{log(k+1)-log k}{tan^{-1}(2/k)}\]

OpenStudy (rational):

try limit test

OpenStudy (rational):

\[\lim\limits_{k\to\infty} \dfrac{log(k+1)-log k}{tan^{-1}(2/k)}= ?\]

OpenStudy (loser66):

I know ln (1+1/k) goes to 0 if k goes to infinitive. How to argue with denominator?

OpenStudy (rational):

lhopital allowed ?

OpenStudy (ikram002p):

Ok with first look,convergence

OpenStudy (loser66):

May be!! I am allowed to use whatever.

OpenStudy (rational):

use lhopital rule then

OpenStudy (loser66):

But it turns out to convergence when the answer from the book is divergence

OpenStudy (rational):

how do you say it converges ?

OpenStudy (ikram002p):

Basically cuz im talking about the one on the question xD

OpenStudy (rational):

yes the first series converges, im asking loser how he is saying the second series with logs converges

OpenStudy (loser66):

numerator: (log (1+1/k))'= 0 denominator (tan^-1 (2/k)' = -2/(k^2+4)

OpenStudy (loser66):

ok, it is divergent. Thank you

OpenStudy (loser66):

Give me the trick, please. How can you quickly figure out which method we should apply?

OpenStudy (rational):

why is it divergent ?

OpenStudy (loser66):

as k --> infinitive, lim is undefined

OpenStudy (rational):

the series diverges by limit test because the limit is NOT 0.

OpenStudy (rational):

Limit test : \(\lim\limits_{n\to\infty}a_n\ne0 ~~\implies ~~~\sum\limits_{n=1}^{\infty} a_n\) diverges

OpenStudy (loser66):

MY God!! I confused everything.

OpenStudy (rational):

thats okay, so basically one always tries "limit test" first, if the limit is not 0, then the series diverges.

OpenStudy (rational):

here is my preferred order of tests for series with positive terms : 1) limit test 2) ratio test 3) comparison test 4) limit comparison test 5) root test 6) others

OpenStudy (loser66):

You mean I have to take one by one, if limit test fail, then try ratio, then....., right?

OpenStudy (rational):

upto you, if limit test is inconclusive, then i generally think of ratio test next but thats not a strict rule, i sometimes try comparison test directly for series like : \(\sum\limits_{n=0}^{\infty}\dfrac{1}{n^2+1}\)

OpenStudy (rational):

i like comparison test more but ratio test gets the work done quickly for most series in textbooks

OpenStudy (loser66):

For comparison test, we need manipulate the original series to get the familiar form, how to quickly figure it out? Like the first one of my problem.

OpenStudy (loser66):

If I use comparison instead of p-series, how can I find \(b_n\)? the series I want to compare to \(a_n\)

OpenStudy (rational):

practice i guess, i don't have any good tips on recognizing when to use which test...

OpenStudy (loser66):

Ok, Thanks a lot. If it is "practice", I have to. No choice, right?

OpenStudy (rational):

the first series is worked using "comparison test", we compared it with the familiar p-series : Since \(\dfrac{\sqrt{n+1}}{n^2-3n+1} \lt\dfrac{1}{(n-4)^{3/2}}\) for \(n\gt 4\), and the series \(\sum\limits_{n=5}^{\infty}\dfrac{1}{(n-4)^{3/2}}\) converges by p-series test, the actual series \(\sum\limits_{n=5}^{\infty}\dfrac{\sqrt{n+1}}{n^2-3n+1} \) converges by "comparison test".

OpenStudy (rational):

keep in mind it is "comparison test", not p-series test. p-series is the series we compared against

OpenStudy (loser66):

What if I use comparison test with \(b_n= \dfrac{1}{(n+1)^{3/2}}\)

OpenStudy (rational):

do you mean "limit comparison test" ?

OpenStudy (rational):

that works equally well!

OpenStudy (loser66):

yes

OpenStudy (loser66):

Thank you so much.

OpenStudy (rational):

np:)